OpenStudy (anonymous):

Donna's company reimburses her expenses on food, lodging, and conveyance during business trips. The company pays $50 a day for food and lodging and $0.60 for each mile traveled. Donna drove 200 miles and was reimbursed $1,620. Part A: Create an equation that will determine the number of days x on the trip Part B: Solve this equation justifying each step with an algebraic property of equality. Part C: How many days did Donna spend on this trip?

OpenStudy (anonymous):

@thehumantorch10

OpenStudy (ksaimouli):

A) \[50x+120=1620\]

OpenStudy (ksaimouli):

120 because 200(0.60)

OpenStudy (ksaimouli):

solve for x for part B,C

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

i just need B amd C

OpenStudy (haseeb96):

as we know the answer of A is 50x +120 = 1620 here x is the number of days . so what we will do here to find part C we will just isolate x it will be 50x = 1620 -120 50x = 1500 x = 1500/50 x = 30 So he will spend 30 days on trip .

OpenStudy (haseeb96):

For part B what we will do 50x +120= 1620 taking -120 on both sides 50x + 120 -120 = 1620 - 120 50x = 1500 now divide both sides by 50, we get x=30 this is your Part B

OpenStudy (haseeb96):

did you get it?

OpenStudy (anonymous):

yes i did thnx :)